# ParaICL: Towards Robust Parallel In-Context Learning

### v1

@@ 4.4 Experiment Results
Retrieval (5-shot) 22.1 28.3
ParaICL (15-shot) 24.9 31.3
@@
5 Ablation Studies and Analysis
5.1 Number of Batches
ParaICL keeps the context length for each batch manageable, allowing for the increasing of batch numbers as hardware capabilities allow. In this experiment, we set the number of demonstration examples in each batch as five, and progressively increase the number of batches. The results of Mistral-7B-Instruct-v0.2 on GSM8K is in Figure 4. With the increment in batch numbers, ParaICL’s improvements tend to converge at five batches. Further increase in batch numbers leads to instability in results. We attribute this to the higher likelihood of incorporating batches that detrimentally affect token selection. Consequently, we set the batch number to three in our main experiments to ensure stable performance improvements.
5.2 Majority Voting vs. Standard Average vs. Weighted Average
Method Llama-2 Mistral
w/ $\mathcal{V}_{head}$ (9-shot) 25.4 42.8
w/o $\mathcal{V}_{head}$ (9-shot) 2.9 3.1
Table 4: Results of ParaICL on GSM8K with and without $\mathcal{V}_{head}$.
@@ 5.1 Number of Batches
5.3 Adaptive plausibility constraint
5.4 Close-Source LLMs
Method GSM8K HellaSwag
Standard (9-shot) 62.0 82.0
ParaICL (9-shot) 66.0 88.0
Table 5: Experimental results of ParaICL on GSM8K and HellaSwag using gpt-3.5-turbo-instruct.
Method Llama-2 Mistral
Standard (9-shot) 21.9 41.5
ParaICL (9-shot) 25.4 42.8
ParaICL w. CD (9-shot) 26.1 43.1
Standard (15-shot) 21.1 40.4
ParaICL (15-shot) 24.9 42.9
ParaICL w. CD (15-shot) 25.2 43.5
Table 6: Results of ParaICL with contrastive decoding on GSM8K.
Our method necessitates the next token probabilities from the model (i.e., the softmax of the raw scores generated by the final layer) for computing the semantic weighted average. However, close-source LLMs, due to their proprietary nature, do not make these probabilities available. For instance, OpenAI’s gpt-3.5-turbo-instruct model only allows access to a maximum of five tokens that have top log probabilities, which represents the extent of information available for our use. We employ the same steps as shown in Section 3.3 to execute parallel semantic decoding using the provided log probabilities from OpenAI models. Concerns regarding API costs lead us to limit our experimentation to a randomly chosen set of 50 data points from the GSM8K and HellaSwag datasets, respectively. Table 6 illustrates that ParaICL enhances performance beyond the standard method, further showcasing its adaptability even with limited information of token distributions.
5.5 Integration with Other Methods

### v2

@@ 5.1 Number of Batches
w/o S.S. (9-shot) 21.1 32.2
w/ S.S. (9-shot) 25.4 33.9
Table 3: With vs. without semantic sorting on GSM8K and HellaSwag. S.S. stands for semantic sorting.
5.2 Parallel Batching with vs. without Semantic Sorting
The demonstration examples are sequenced by their question semantic similarities to the test question before parallel batching. This ensures that semantically similar examples are batched together, allowing them to share the same batch similarity score during decoding. According to results shown in Table 3, parallel batching with semantic sorting significantly outperforms without semantic sorting. In fact, Parallel batching without semantic sorting even performs worse than the standard approach on GSM8K. This is because, without semantic sorting, a single batch can contain both relevant and irrelevant demonstration examples. This mix of examples reduces the overall decoding probability from the relevant examples in the batch.
Method Llama-2 Mistral
w/ $\mathcal{V}_{head}$ (9-shot) 25.4 42.8
w/o $\mathcal{V}_{head}$ (9-shot) 2.9 3.1
Table 4: Results of ParaICL on GSM8K with and without $\mathcal{V}_{head}$.
Method Llama-2-13B-Chat Llama-3.1-70B-Instruct Mixtral-8x7B-Instruct-v0.1 gpt-3.5-turbo-instruct
Standard (9-shot) 23.1 86.4 58.1 62.0
ParaICL (9-shot) 25.6 87.1 59.2 66.0
Table 5: Experimental results of ParaICL on GSM8K using more open- and closed-source LLMs.
Method Llama-2 Mistral
Standard (9-shot) 21.9 41.5
ParaICL (9-shot) 25.4 42.8
ParaICL w. C.D. (9-shot) 26.1 43.1
Standard (15-shot) 21.1 40.4
ParaICL (15-shot) 24.9 42.9
ParaICL w. C.D. (15-shot) 25.2 43.5
Table 6: Results of ParaICL with contrastive decoding on GSM8K. C.D. stands for contrastive decoding.
5.3 Adaptive Plausibility Constraint
5.4 More Open- and Closed-Source LLMs
In this section, we demonstrate that ParaICL is effective in both open- and closed-source LLMs.
We first conduct experiments on GSM8K using three additional open-source models spanning a range of parameter counts, including Llama-2-13B-Chat, Llama-3.1-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1. As shown in Table 5, ParaICL consistently enhances the performances, highlighting its adaptability to open-source models.
Our method necessitates the next token probabilities from the model (i.e., the softmax of the raw scores generated by the final layer) for computing the semantic weighted average. However, closed-source LLMs, due to their proprietary nature, do not make these probabilities available. For instance, OpenAI’s gpt-3.5-turbo-instruct model only allows access to a maximum of five tokens that have top log probabilities, which represents the extent of information available for our use. We employ the same steps as shown in Section 3.3 to execute parallel semantic decoding using the provided log probabilities from OpenAI models. Concerns regarding API costs lead us to limit our experimentation to a randomly chosen set of 50 data points from the GSM8K datasets. Table 5 illustrates that ParaICL enhances performance beyond the standard method, further showcasing its adaptability even with limited information on token distributions. In contrast, PCW is restricted to open-source models as it requires modifications to the attention mask.
5.5 Integration with Other Methods
5.6 Majority Voting vs. Standard Average vs. Weighted Average
@@ 5.5 Integration with Other Methods
5.7 Performance-Efficiency Tradeoff
ParaICL introduces additional inference time due to the computation of embeddings for both the demonstrations and test examples. Nevertheless, this overhead is minimal when computations are conducted in parallel. Across various datasets, the average time to compute embeddings using SimCSE (Gao et al., 2021) is merely 5.1 seconds.
@@ 6 Conclusions
Limitations
We only study the effectiveness of ParaICL in decoder-only language models. In future works, we aim to extend our evaluation of ParaICL to language models that utilize different structures, including those with encoder-only and encoder-decoder structures. In this paper, we only demonstrate that ParaICL can be easily integrated with contrastive objectives. We would like to explore the integration of ParaICL with other methods, such as retrieval-augmented generation.
Ethical Impact
We do not foresee any potential ethical issues with our proposed method.
Acknowledgements
This work was supported by DAMO Academy through DAMO Academy Research Intern Program.